### v1

# LeGrad: An Explainability Method for Vision Transformers via Feature Formation Sensitivity

Abstract
Vision Transformers (ViTs), with their ability to model long-range dependencies through self-attention mechanisms, have become a standard architecture in computer vision. However, the interpretability of these models remains a challenge. To address this, we propose LeGrad, an explainability method specifically designed for ViTs. LeGrad computes the gradient with respect to the attention maps of ViT layers, considering the gradient itself as the explainability signal.We aggregate the signal over all layers, combining the activations of the last as well as intermediate tokens to produce the merged explainability map. This makes LeGrad a conceptually simple and an easy-to-implement tool for enhancing the transparency of ViTs. We evaluate LeGrad in challenging segmentation, perturbation, and open-vocabulary settings, showcasing its versatility compared to other SotA explainability methods demonstrating its superior spatial fidelity and robustness to perturbations. [^1] A demo and the code is available at https://github.com/WalBouss/LeGrad
Figure 1: LeGrad explainability maps: For a given vision-language model and a textual prompt, LeGrad generates a heatmap indicating the part of the image that is most sensitive to that prompt. LeGrad performs equally well across a large range of model sizes, ranging from ViT-B/16(150M params.) to ViT-bigG/14(2B params.).
1 Introduction
In the realm of computer vision, Vision Transformers (ViTs)[16] have revolutionized the field with their ability to model long-range dependencies through self-attention mechanisms [43, 27, 45, 4, 30]. Explanability methods designed for convolutional or feed-forward neural networks are not directly applicable to ViTs due to their architectural requirements, like GradCAM’s [37] reliance on convolutional layers and Layer-wise Relevance Propagation’s (LRP) [5] specific layer-wise propagation rules. While ViT-specific interpretability techniques exist, including adaptations of traditional methods [12, 37, 13], attention-based techniques [1, 44, 12, 13], and text-based explanations [22, 20, 1], the interpretability especially of large-scale architectures remains a challenge.
To address this, we propose a Layerwise Explainability method that considers the Gradient with respect to the attention maps, LeGrad. LeGrad is specifically designed for ViTs as leverages the self-attention mechanism intrinsic to ViTs to generate relevancy maps highlighting the most influential parts of an image for the model’s prediction. Compared to other methods, LeGrad computes gradient with respect to the attention maps, as e.g. opposed to CheferCam [12, 13], which uses the gradient to weight the attention maps. This is done independently for each layer. The final explainability signal is then pooled over all layers of the ViT. Note that using a layerweise gradient, compared to other signals, allows to sum up over different layers without further need for normalization. To further improve the signal, the gradient is clipped by a ReLU function preventing negative gradients to impact positive activations. The proposed approach is conceptually simple and versatile, as it only requires the gradient w.r.t. to the ViT’s attention maps. This facilitates its adoption across various applications and architecture, including larger ViTs such as ViT-BigG as well as attention pooling architectures [26, 49], enhancing the transparency of ViTs.
We evaluate the proposed method for various ViT backbones on three challenging tasks: segmentation, open-vocabulary detection, and perturbation on ImageNet [34, 19] as well as OpenImagesV7 [6]. It shows that while current methods struggle especially with the diverse object categories in OpenImagesV7, LeGrad can achieve performance gains of 2×2$\times$2 ×5×5$\times$5 ×, reaching a score of $48.4$ p-mIoU on OpenImagesV7 using OpenCLIP-ViT-B/16. Furthermore, we demonstrate the applicability of LeGrad to very large models, such as the ViT-BigG/14 [14] with 2.5 billion parameters. Finally, it also adapts well to different feature aggregation strategies employed by ViTs [32, 49], making it a versatile tool for explainability.
We summarize the contributions as follows: (1) We propose LeGrad as a layerwise explainability method based on the gradient with respect to ViTs attention maps. (2) As the layerwise explainability allows to easily pool over many layers, LeGrad scales to large architectures such as ViT-BigG/14 and is applicable to various feature aggregation methods. (3) We evaluate LeGrad on various tasks and benchmarks, showing its improvement compared to other state-of-the-art explainability methods especially for large-scale open vocabulary settings.
2 Related Work
Gradient-Based Explanation Methods Feature-attribution methods (also known as saliency methods) are a commonly used explanation technique that explain model decisions by assigning a score to each image pixel representing its importance to the model’s output. Largely, these methods can be categorized into two groups [29] — gradient-based methods that compute explanations based on the gradient of the model’s prediction with respect to each input pixel [38, 17, 41, 40, 39, 24, 37] and perturbation-based methods that measure pixel-importance by successively perturbing the input images and measuring the impact on the model’s output [28, 33, 31, 11, 48]. While both types of methods have been used successfully to identify reasoning, spurious correlation, and trustworthiness in traditional computer vision models [7, 10], gradient-based methods are often more computationally efficient since they only require a single backwards pass and simple to understand since they are a direct function of the model’s parameters and do not rely on additional models or image modifications. However, many existing gradient-based methods were designed for convolutional and feed-forward model architectures, so it is non-trivial to directly apply them to ViTs since ViTs do not contain spatial feature maps and include complex interactions between patches induced by the self-attention mechanism. Instead, we build on the simplicity and computational efficiency of gradient-based explanation methods to develop LeGrad, a gradient-based feature-attribution method specifically designed for ViT architectures.
While most gradient-based methods were designed prior to the widespread use of ViTs, researchers have recently made efforts to adapt existing methods or develop new ones specifically for transformers. Chefer et al. [12] extend LRP [5] to transformers by integrating gradients within the self-attention layers. However, this approach is computationally heavy and is inflexible to architecture changes as it requires specific implementation for each module of the network. To circumvent that complexity, CheferCAM [13] weights the attention by their gradient and aggregates it through the layer via matrix multiplication. However, the use of gradients to weigh the attention heads’ importance makes this method class-specific. LeGrad on the other hand is class-agnostic and computationally efficient, making it simple to use and more effective than these other methods (see Sec. 4 for a quantitative evaluation).
Attention-Based Explanation Methods A separate line of research has proposed using attention maps, as opposed to gradients, as a way to explain for transformers’ decisions [1, 44]. One attention-based method, rollout [1], traces the flow of importance through the transformer’s layers by linearly combining the attention maps via matrix multiplication. Attention flow [1] contextualizes the attention mechanism as a max-flow problem; however, it is computationally demanding and has not been extensively evaluated for vision tasks. While these methods offer insights into the attention mechanism, they often neglect the non-linear interactions between attention heads and the subsequent layers. Moreover, they may not adequately distinguish between positive and negative contributions to the final decision, leading to potentially misleading interpretations, as found in Chefer et al. [12]. While LeGrad also focuses on the attention maps to generate its explanation, it uses the gradient w.r.t. to the attention maps, thereby assessing the sensitivity of the attention maps to a change in the patch tokens. LeGrad does not attempt to track the flow of information throughout the transformer layers but, rather, focuses on assessing the sensitivity of the model prediction at each layer for every location of the image.
Text-based Explanations of Intermediate Representations Research has also explored text-based methods for interpreting intermediate representations in vision models. For instance, leveraging CLIP’s language-image space, researchers have provided text descriptions for active neuron regions [22, 20] and projected model features into text-based concept banks [18]. In particular, TextSpan [18] focuses on the explainability of CLIP-like models. It refrains from using gradient computation by aggregating the intermediate features’ similarities along a given text direction, creating a relevancy map for a text query. LeGrad advances this line of work by focusing on the sensitivity of feature representations within ViTs, offering a clear and concise method to generate relevancy maps that can be adapted to various feature aggregation strategies.
Evaluating Explanation Methods As the popularity of feature-attribution based explanation methods has grown, so has research evaluating their ability to faithfully reflect the model’s decision making process [41, 46, 25, 35, 23, 3, 2, 42, 21, 50]. Faithfulness is an important criteria because explanation methods that highlight human-salient pixels may look correct even if the model relies on spurious correlations and give human interpreters unwarranted confidence in the model. These evaluations can be categorized based on the aspect of faithfulness that they test [8], such as the method’s sensitivity to meaningful changes to the input, model, or data [41, 46, 25, 35, 23, 3, 2, 42] and its ability to highlight a succinct and meaningful pixel region [21, 50]. We evaluate LeGrad on both of these types of tests, demonstrating that it is faithful to the model’s decision making process by testing that removing LeGrad’s important pixels impacts the model’s output and checking that LeGrad’s output corresponds to a concise object region.
Figure 2: Overview of LeGrad: Given a text prompt or a classifier $\mathcal{C}$, for each layer $l$, an activation $s^{l}$ is computed and used to compute the explainability of that layer. The layerwise explainability maps are then merged to produce LeGrad’s output.
3 Method
In this section, we first introduce ViT’s mechanics and the different feature aggregation mechanisms used for this architecture. We then explain the details of LeGrad, starting by a single layer and then extending it to multiple layers.
3.1 Feature formation in ViTs
The ViT architecture is a sequence-based model that processes images by dividing them into a grid of $n$ patches. These patches are linearly embedded and concatenated with a class token $z_{0}^{0}=z^{0}_{[CLS]}\in\mathbb{R}^{d}$, which is designed to capture the global image representation for classification tasks. The input image $I$ is thus represented as a sequence of $n+1$ tokens $Z^{0}=\{z_{0}^{0},z_{1}^{0},\dots,z_{n}^{0}\}$, each of dimension $d$, with positional encodings added to retain spatial information.
The transformation of the initial sequence $Z^{0}\in\mathbb{R}^{(n+1)\times d}$ through the ViT involves $L$ layers, each performing a series of operations. Specifically, each layer $l$ applies multi-head self-attention (MSA) followed by a multilayer perceptron (MLP) block, both with residual connections:
$\hat{Z}^{l}$ $=\text{MSA}^{l}(Z^{l-1})+Z^{l-1},$ (1)
$Z^{l}$ $=\text{MLP}^{l}(\hat{Z}^{l})+\hat{Z}^{l}.$
After $L$ layers, the image representation can be obtained via various strategies:
[CLS] token: The class token approach, as introduced in the original ViT paper [16], uses the processed class token as the image embedding $\bar{z}_{[CLS]}=z^{L}_{0}$. This method relies on the transformer’s ability to aggregate information from the patch tokens into the class token during training.
Attentional Pooler: as used in [26, 47, 49], employs an additional multi-head attention layer with a learnable query token $q_{pool}\in\mathbb{R}^{d}$. This token interacts with the final layer patch tokens to produce the pooled representation $\bar{z}_{AttnPool}$:
$\bar{z}_{AttnPool}=\text{softmax}\left(\frac{q_{pool}\cdot(W_{K}Z^{L})^{T}}{\sqrt{d}}\right)(W_{V}Z^{L}),$ (2)
where $W_{K},W_{V}\in\mathbb{R}^{d\times d}$ are learnable projection matrices.
Independent of the feature aggregation strategy, it is important for an explainability method to account for the iterative nature of feature formation in ViTs. The goal should be to capture the contributions of all layers for the final representation, providing insights into the model’s decision-making process.
3.2 Explainability Method: LeGrad
We denote the output tokens of each block $l$ of the ViT as $Z^{l}=\{z_{0}^{l},z_{1}^{l},\dots,z_{n}^{l}\}\in\mathbb{R}^{(n+1)\times d}$, where $d$ is the dimensionality of each token.
Consider a classifier $\mathcal{C}\in\mathbb{R}^{d\times C}$, where $C$ is the number of classes. This classifier can be learned during training for a supervised classification task (e.g., ImageNet) or can be formed from text embeddings of class descriptions for a zero-shot classifier, as in the case of vision-language models like CLIP, generating a prediction $\bar{y}$. This prediction $\bar{y}$ is obtained by passing the aggregated feature representation of the ViT, noted $\bar{z}$, through the classifier $\mathcal{C}$ and selecting the class with the highest activation:
$\bar{y}=\bar{z}\cdot\mathcal{C}\in\mathbb{R}^{C},$ (3)
The following is the description of how to obtain the explainability map for a single layer, using the last layer as an example and then how to generalize it to multiple layers. The overall method is visualized Figure 2.
Figure 3: LeGrad for a single layer.
Process for a Single Layer: To compute a 2D map that highlights the image regions most influential for the model’s prediction of a particular class, we focus on the activation with respect to the target class $\hat{c}$, denoted by $s=\bar{y}_{[\hat{c}]}$. The attention operation within a ViT is key to information sharing, and thus our method concentrates on this process.
We compute the gradient of the activation $s$ with respect to the attention map of the last layer, as shown in Figure 3, denoted as $\mathbf{A}^{L}\in\mathbb{R}^{h\times n\times n}$:
$\nabla\mathbf{A}^{L}=\frac{\partial s}{\partial\mathbf{A}^{L}}\in\mathbb{R}^{h\times(n+1)\times(n+1)},$ (4)
where $h$ is the number of heads in the self-attention operation. Negative contributions are eliminated by applying a ReLU function, and the gradient is averaged across the patch and head dimensions:
$\hat{E^{L}}(s)=\frac{1}{h\cdot(n+1)}\sum_{h}\sum_{i}\left(\nabla\mathbf{A}^{L}_{h,i,.}\right)^{+}\in\mathbb{R}^{n+1}.$ (5)
The final explainability map is obtained by removing the column corresponding to the [CLS] token to isolate the influence of the patch tokens, reshaping it into a 2D map, and applying min-max normalization:
$E^{L}(s)=\text{norm}(\text{reshape}(\hat{E}^{L}(s)_{1:}))\in\mathbb{R}^{W\times H}.$ (6)
Process for Multiple Layers: Recognizing that information aggregation occurs over several layers, we extend the process to all layers. For each layer $l$, we calculate the activation score $s^{l}$ using the intermediate tokens $Z^{l}$ and derive the explainability map accordingly:
$\hat{E}^{l}(s^{l})=\frac{1}{h\cdot(n+1)}\sum_{h}\sum_{i}\left(\nabla\mathbf{A}^{l}_{h,i,.}\right)^{+}\in\mathbb{R}^{n+1}.$ (7)
Prior to normalization, we average the explainability maps from each layer:
$\bar{\mathbf{E}}$ $=\frac{1}{L}\sum_{l}\hat{E}^{l}(s^{l})_{1:}$ (8)
$\mathbf{E}$ $=\text{norm}(\text{reshape}(\bar{\mathbf{E}}))\in R^{W\times H}.$
This approach queries each patch token at a given layer about its influence on the prediction at that stage.
Adaptation to Attentional Pooler: For ViTs using an attentional pooler (e.g. SigLIP [49]), a slight modification is made to compute the activation $s^{l}$ at each layer. We apply the attentional pooler module $Attn_{pool}$ to each intermediate representation $Z^{l}$ to obtain a pooled query $q^{l}\in\mathbb{R}^{d}$. The activation $s^{l}$ with respect to the desired class $c$ is then computed as $s^{l}=q^{l}\cdot\mathcal{C}_{:,c}\in\mathbb{R}$. Instead of considering the self-attention map, we use the attention map of the attentional pooler, denoted $\mathbf{A}_{pool}\in\mathbb{R}^{h\times 1\times n}$. Thus, for every layer $l$, $\nabla A^{l}=\frac{\partial s^{l}}{\partial A^{l}_{pool}}$.
4 Experiments
4.1 Evaluation Settings
Implementation details: For the empirical evaluation of LeGrad, we utilized a suite of Vision Transformer models sourced from the OpenCLIP library [14], trained on the LAION dataset [36]. The models employed in our experiments include ViT-B/16, ViT-L/14, ViT-H/14, and ViT-bigG/14.For SigLIP [49], we use the ViT-B/16 with the "webli" weights. Each model processes input images at a resolution of $224\times 224$ pixels. A key advantage of LeGrad is its hyperparameter-free nature, which obviates the need for hyperparameter tuning. Furthermore, we do not retraining or fine-tune the vision-language models, thereby demonstrating their capabilities to localize queries by solely being trained on image-level data without any localization annotations during training or fine-tuning phases.
[…]
4.2 Comparison to State-of-the-Art
Segmentation Analysis: We first evaluate LeGrad as well as other methods in the context of image segmentation using the ImageNet-segmentation dataset. As shown in Table 2, LeGrad achieved a mIoU of 58.7%, surpassing other SOTA explainability methods. Notably, it outperformed CheferCAM as a gradient-based method for ViTs, and TextSpan as a non-gradient-based method, indicating its robustness in capturing relevant image features for classification tasks.
Open-Vocabulary Performance: We further evaluate the performance on the OpenImagesV7 dataset testing the capabilities of all methods in handling diverse object categories. Note that for GradCAM we searched over the layers and took the one that was performing the best. As shown in Table 2, LeGrad outperforms all other SOTA methods, with performance gains ranging from 2×2$\times$2 × to 5×5$\times$5 × compared to the second-best performing method. This can be seen as an indicator for LeGrad’s capacity for fine-grained recognition.
Perturbation-Based Evaluation: We further performed the perturbation-based evaluation on the ImageNet-val datasets. As detailed in Table 3, LeGrad’s performance is here comparable to TextSpan for positive perturbations and slightly superior for negative perturbations. For all other methods, LeGrad outperformes both attention-based (e.g., "rollout" and "raw attention") and gradient-based methods (e.g., GradCAM, AttentionCAM, and CheferCAM) across various model sizes and for both predicted and ground truth classes, emphasizing its ability to identify and preserve critical image regions for accurate classification.
Performance on SigLIP: We finally evaluate the adaptability b regarding the performance on SigLIP-B/16, a Vision-Language model employing an attentional pooler as shown in Table 4. The results underscore the methods performance across both negative and positive perturbation-based benchmarks. Notably, in the open-vocabulary benchmark on OpenImagesV7, LeGrad achieved a p-mIoU of 25.4, significantly surpassing GradCAM’s 7.0 p-mIoU, the next best method. These findings affirm the versatility of LeGrad, demonstrating its robust applicability to various pooling mechanisms within Vision Transformers. Further details on the methodological adaptations of LeGrad and other evaluated methods for compatibility with SigLIP are provided in the annex.
[…]
LRP 52.81 33.57 54.37
Partial-LRP 61.49 40.71 72.29
rollout 60.63 40.64 74.47
Raw attention 65.67 43.83 76.05
GradCAM 70.27 44.50 70.30
CheferCAM 69.21 47.47 78.29
TextSpan 73.01 40.26 81.4
LeGrad 77.52 58.66 82.49
Table 1: Object Segmentation: method comparison on ImageNet-S using a CLIP(ViT-B/16) model.
p-mIoU $\uparrow$
Method B/16 L/14 H/14 bigG/14
rollout 8.75 6.85 5.82 4.74
Raw attention 0.94 1.60 0.85 0.07
GradCAM 8.72 2.80 2.46 0.14
AttentionCAM 5.87 4.74 1.20 0.72
CheferCAM 5.87 2.51 9.49 9.71
TextSpan 9.44 21.73 23.74 20.18
LeGrad 48.38 47.69 46.51 36.47
Table 2: Open-Vocabulary Segmentation Performance: methods comparison on OpenImagesV7 using different openCLIP model size.
Negative Positive
Method Predicted $\uparrow$ Target $\uparrow$ Predicted $\downarrow$ Target $\downarrow$
ViT-B/16 rollout 44.36 44.36 23.03 23.03
Raw attention 46.97 46.97 20.23 20.23
GradCAM 32.01 45.26 36.52 22.86
AttentionCAM 39.56 39.68 34.28 34.12
CheferCAM 47.91 49.28 18.66 17.70
TextSpan 50.92 52.81 15.10 14.26
LeGrad 50.24 52.27 15.06 13.97
ViT-L/14 rollout 40.46 40.46 29.46 29.46
Raw attention 47.14 47.14 23.49 23.49
GradCAM 45.24 47.08 23.81 22.68
AttentionCAM 45.81 45.84 31.18 31.03
CheferCAM 49.69 50.37 20.67 20.14
TextSpan 53.17 54.42 16.77 16.12
LeGrad 53.11 54.48 15.98 15.23
Negative Positive
Method Predicted $\uparrow$ Target $\uparrow$ Predicted $\downarrow$ Target $\downarrow$
ViT-H/14 rollout 49.37 49.37 32.91 32.91
Raw attention 54.42 54.42 29.25 29.25
GradCAM 45.26 45.48 40.98 40.91
AttentionCAM 51.62 51.65 36.72 36.56
CheferCAM 56.55 56.56 26.17 26.16
TextSpan 60.14 61.91 20.07 19.14
LeGrad 60.02 61.72 19.30 18.26
ViT-BigG/14 rollout 38.44 38.44 48.72 48.72
Raw attention 56.56 56.56 31.25 31.25
GradCAM 40.06 40.98 57.53 56.31
AttentionCAM 54.68 54.74 39.28 39.27
CheferCAM 57.95 58.27 29.21 28.45
TextSpan 63.32 64.95 21.96 21.06
LeGrad 62.62 64.67 21.33 20.28
Table 3: SOTA Perturbation Performance: Comparison of interpretability methods on the ImageNet-val using a different model size.
ImageNet OpenImagesV7
Negative Positive
Method Predicted $\uparrow$ Target $\uparrow$ Predicted $\downarrow$ Target $\downarrow$ p-mIoU $\uparrow$
rollout 47.81 47.81 25.74 25.74 0.07
Raw attention 44.42 44.42 25.85 25.85 0.09
GradCAM 41.25 44.42 35.10 33.50 6.97
AttentionCAM 45.62 45.71 45.01 44.92 0.19
CheferCAM 47.12 49.13 22.35 21.15 1.94
LeGrad 50.08 51.67 18.48 17.55 25.40
Table 4: SOTA comparison on SigLIP-B/16: Comparison of interpretability methods on perturbation-based tasks on ImageNet-val and open-vocabulary localization on OpenImagesV7.
4.3 Ablation
4.3.1 Ablation Studies on Layer Contributions
Ablation studies are conducted to understand the impact of the number of layers considered in the computation of LeGrad’s explainability maps. We investigate this aspect across two distinct benchmarks: a perturbation-based evaluation using the ImageNet validation set and an open-vocabulary segmentation task on the OpenImagesV7 dataset. The experiments are performed on various model sizes from the openCLIP library, including ViT-B/16, ViT-L/14, and ViT-H/14.
Perturbation-Based Evaluation: In the perturbation-based evaluation, see Figure 4(a), we employ a negative perturbation test as described in Section 4.1, using the ground-truth class for reference. The results indicate that the ViT-B/16 model’s performance is optimal when fewer layers are included in the explainability map computation. Conversely, larger models such as ViT-L/14 and ViT-H/14 show improved performance with the inclusion of more layers. This suggests that in larger models, the aggregation of information into the [CLS] token is distributed across a greater number of layers, necessitating a more comprehensive layer-wise analysis for accurate explainability.
Open-Vocabulary Segmentation: For the open-vocabulary segmentation task, Figure 4(b), all models demonstrate enhanced performance with the inclusion of additional layers in the explainability map computation. The optimal number of layers varies with the model size, with larger models requiring a greater number of layers. This finding aligns with existing literature [18], which suggests that the information aggregation process in ViTs is more distributed in larger architectures. However, it is also observed that beyond a certain number of layers, the performance gains plateau, indicating that the inclusion of additional layers does not further enhance the explainability map’s quality.
These ablation studies underscore the importance of layer integration in the construction of explainability maps for Vision Transformers. They further show that while the inclusion of more layers can be beneficial, especially for larger models, there exists an optimal range beyond which no additional interpretability is gained. Indeed, we note in Figure 4 that including the last $40\%$ of the ViT layers in LeGrad’s explainability computation is a good trade-off. This can be helpful for the practical application of LeGrad, as it allows for a more targeted and computationally efficient approach to generating explainability maps without compromising on their quality.
(a) AUC for Negative perturbation on ImageNet-val for different layers used for LeGrad
(b) point-mIoU on OpenImagesV7 for different layers used for LeGrad.
Figure 4: Ablation on the number of layers used in LeGrad for different architecture sizes.
4.3.2 Ablations of ReLU and Layer Aggregation
We finally scrutinize the design choices underpinning LeGrad in Table 5. Specifically, we investigate the effect of discarding negative gradients before aggregating layer-specific explainability maps vai ReLU, as well as the implications of leveraging intermediate feature tokens $Z^{l}$ to compute gradients for each respective layer. We use the framework of the perturbation benchmark delineated in Section 4.1 and both ViT-L/14 and ViT-H/14 models. The results indicate that the omission of either component induces decline in performance, thereby affirming the role these elements play in the architecture of the method.
L/14 H/14
[…]
47.81 20.80 57.57 21.73
49.32 19.95 59.55 19.50
52.01 16.80 60.28 18.26
54.48 15.23 61.72 18.26
Table 5: Ablation study: "ReLU" corresponds to whether or not negative gradients are set to 0. "All layers" corresponds to whether or not the intermediate tokens are used to compute the gradient for every layer or if only the features from the last layer are used. Numbers are the AUC score for the perturbation base benchmark using the target class to compute the explainability map.
4.4 Qualitative analysis
4.4.1 Ablation Studies on Layer Contributions
To elucidate the interpretability of Vision Transformers (ViTs) through our proposed LeGrad method, we extend our examination beyond quantitative ablation studies to a qualitative analysis. This analysis scrutinizes the localization maps generated by LeGrad for the last ten layers of various ViT models, as well as the final map that merge these individual layer contributions. The experiments are performed on various model sizes from the openCLIP library, including ViT-B/16, ViT-L/14, and ViT-H/14.
The qualitative results, illustrated in Figure 5, reveal that the localization of the prompt is not confined to a single layer but is distributed across multiple layers. This observation is consistent with the findings from the ablation studies in Section 4.3, which underscore the necessity of incorporating multiple layers into the explainability framework to capture the full scope of the model’s decision-making process.
A notable distinction is visible in the behavior of smaller models, such as ViT-B/16, with that of larger counterparts like ViT-L/14 and ViT-H/14. In the case of ViT-B/16, the explainability signal predominantly emanates from the final layers, particularly the last two. Conversely, in larger models, earlier layers also substantially influence the merged explainability map. For instance, in Figure 5, layers 8 and 9 of ViT-H/14 appear to distinctly focus on the heads of the cats, while other layers’ contributions are more diffused. This observation not only corroborates the utility of incorporating multiple layers into the explainability analysis but also suggests a more distributed information aggregation process into the [CLS] token in larger models, as posited in the literature [18].
The qualitative analysis thus reinforces the hypothesis that the mechanism by which ViTs integrate information across layers is size-dependent, with larger models engaging a broader range of layers in the synthesis of the global feature representation. This is important for the practical application of LeGrad, as it informs the selection of layers to be included in the explainability map computation, ensuring both computational efficiency and the retention of fidelity.
Figure 5: Qualitative analysis of the impact of each layer for different model sizes using "a photo of a cat" as prompt. In smaller models the explainability signal predominantly emanates from the final layers while in larger models lower layers also contribute to the explainability map.
4.4.2 Qualitative Comparison to SOTA
Here, we present a qualitative analysis of the explainability maps generated by LeGrad in comparison to other state-of-the-art (SOTA) methods. The visual results are depicted in Figure 6, which includes a diverse set of explainability approaches such as gradient-based methods (e.g., CheferCAM, GradCAM), attention-based methods (e.g., Raw Attention weights visualization, Rollout), and methods that integrate intermediate visual representations with text prompts (e.g., TextSpan).
Our observations indicate that raw attention visualizations tend to highlight a few specific pixels with high intensity, often associated with the background rather than the object of interest. This pattern, consistent with findings in the literature [9, 15], suggests that certain tokens disproportionately capture attention weights. Consequently, methods that rely on raw attention weights to construct explainability maps, such as CheferCAM, exhibit similar artifacts. For instance, in the localization of "Wheel" (Figure 6, row 2), the wheel is marginally accentuated amidst a predominance of noisy, irrelevant locations.
In contrast, LeGrad is resilient to such distractions by leveraging gradients with respect to attention maps. The presence of uniform noisy activations across different prompts results in minimal gradients for these regions, effectively filtering them out from the final heatmaps. This characteristic enables LeGrad to produce more focused and relevant visual explanations.
The presence of noisy activations is also observed in methods like TextSpan and GradCAM, which utilize intermediate feature representations. These artifacts tend to surface in the same spatial regions as those highlighted by raw attention weights, suggesting a potential link to tokens with inflated attention. While the underlying cause of this behavior demands further investigation, it is beyond the scope of the current work. Eventually, Figure 6 showcases LeGrad’s capability to accurately localize a wide range of objects and concepts, from commonplace items like "baskets" to abstract notions such as "numbers" and activities like "swimming." The method’s robustness and versatility underscore its effectiveness in providing interpretable insights into the decision-making processes of vision transformers.
Figure 6: SOTA Qualitative Comparison: visual comparison of different explainability methods on images from OpenImagesV7.
5 Conclusion
In this work, we proposed LeGrad as an explainability method that highlights the decision-making process of Vision Transformers (ViTs) across all layers various feature aggregation strategies. We validated the method’s effectiveness in generating interpretable visual explanations that align with the model’s reasoning. Our approach offers a granular view of feature formation sensitivity, providing insights into the contribution of individual layers and attention mechanisms. The evaluation across object segmentation, perturbation-based metrics, and open-vocabulary scenarios underscores LeGrad’s versatility and fidelity in different contexts. By facilitating a deeper understanding of ViTs LeGrad paves the way for more transparent and interpretable foundation models.
Supplementary material
LeGrad: An Explainability Method for Vision Transformers via Feature Formation Sensitivity
7 Overview
In this supplementary material, we first provide in section 8 a link to a demo of our novel method, LeGrad, showcasing its practical application. Detailed information regarding the implementation, including pre-trained weights and baseline methodologies for comparative evaluation, is presented in section 9. An expanded evaluation of LeGrad’s performance, specifically on Vision Transformers (ViT) trained with the ImageNet dataset, is documented in section 10. Section 10.2 offers a visual representation of the perturbation benchmarks utilized to assess the efficacy of various explainability approaches. Additional qualitative examples are provided in section 10.3 to further illustrate the capabilities of our method. Lastly, section 11 includes a disclaimer addressing the use of personal and human subjects data within our research.
8 Demo
An image is worth a $1000$ words, we, therefore, provide a demo at:
https://huggingface.co/spaces/WalidBouss/LeGrad-Demo.
9 Implementation details
9.1 Pretrained weights
The experiments conducted in our study leverage a suite of models with varying capacities, including ViT-B/16, ViT-L/14, ViT-H/14, and ViT-bigG/14. These models are initialized with pretrained weights from the OpenCLIP library respectively identified by: "laion2b_s34b_b88k", "laion2b_s32b_b82k", "laion2b_s32b_b79k", and "laion2b_s39b_b160k". For the SigLIP method, we utilize the ViT-B/16 model equipped with the "webl" weights.
9.2 Detailed Description of Baselines
In this section, we provide a concise overview of the baseline methods against which our proposed approach is compared.
GradCAM: While originally designed for convolutional neural networks (CNNs), GradCAM can be adapted for Vision Transformers (ViTs) by treating the tokens as activations. To compute the GradCAM explainability map for a given activation $s$, we calculate the gradient of $s$ with respect to the token dimensions. The gradients are aggregated across all tokens and serve as weights to quantify the contribution of each token dimension. Formally, for intermediate tokens $Z^{l}=\{z_{0}^{l},z_{1}^{l},\dots,z_{n}^{l}\}\in\mathbb{R}^{(n+1)\times d}$, the GradCAM map $E_{GradCAM}$ is defined as:
$w$ $=\frac{1}{n}\sum_{i=0}^{n}\frac{\partial s}{\partial z_{i}^{l}}\quad\in\mathbb{R}^{d\times 1\times 1}$ (9)
$\hat{E}_{GradCAM}$ $=\left(\frac{1}{d}\sum_{k=1}^{d}w_{d}*Z^{l}_{1:,d}\right)^{+}\quad\in\mathbb{R}^{n}$
$E_{GradCAM}$ $=\text{norm}(\text{resize}(\hat{E}_{GradCAM}))\quad\in\mathbb{R}^{W\times H},$
with $d$ representing the token dimension, * denoting element-wise multiplication, and the superscript $+$ indicating the ReLU operation. We empirically determined that applying GradCAM to layer 8 of ViT-B/16 yields optimal results.
AttentionCAM: This method extends the principles of GradCAM to ViTs by utilizing the attention mechanism within the transformer’s architecture. AttentionCAM leverages the gradient signal to weight the attention maps in the self-attention layers. Specifically, for the last block’s self-attention maps $\mathbf{A}^{L}$, the AttentionCAM map $E_{AttnCAM}$ is computed as:
$\nabla\mathbf{A}^{L}&=\frac{\partial s}{\partial\mathbf{A}^{L}}\in\mathbb{R}^{h\times(n+1)\times(n+1)}\\ w&=\frac{1}{n^{2}}\sum_{i,j}\nabla\mathbf{A}^{L}_{:,i,j}\in\mathbb{R}^{h}\\ \hat{E}_{AttnCAM}&=\sum_{p}^{h}\left(w_{p}*A^{L}_{p,:,:}\right)\in\mathbb{R}^{(n+1)\times(n+1)}\\ E_{AttnCAM}&=\text{norm}(\text{resize}(\hat{E}_{AttnCAM})_{0,1:})\end{aligned},$ (10)
where $h$ denotes the number of heads in the self-attention mechanism.
Raw Attention: This baseline considers the attention maps from the last layer, focusing on the weights associated with the [CLS] token. The attention heads are averaged and the resulting explainability map is normalized. The Raw Attention map $E_{Attn}$ is formalized as:
$\hat{E}_{Attn}$ $=\mathbf{A}^{L}_{:,0,1:}\in\mathbb{R}^{h\times 1\times n}$ (11)
$E_{Attn}$ $=\text{norm}(\text{resize}(\frac{1}{h}\sum_{k=1}^{h}(\hat{E}_{Attn})_{k}))\quad\in\mathbb{R}^{W\times H}$
These baselines provide a comprehensive set of comparative measures to evaluate the efficacy of our proposed method in the context of explainability for ViTs.
9.3 Adaptation of Baseline Methods to Attentional Pooler
In the main manuscript, we introduced our novel method, LeGrad, and its application to Vision Transformers (ViTs) with attentional poolers. Here, we provide supplementary details on how LeGrad and other baseline methods were adapted ViTs employing attentional poolers:
CheferCAM: Following the original paper [13] that introduces CheferCAM, we considered the attentional pooler as an instance of a "decoder transformer" and applied the relevancy update rules described in equation (10) of that paper [13], (following the example of DETR). Since the attentional pooler has no skip connection we adapted the relevancy update rule not to consider the skip connection.
AttentionCAM: For AttentionCAM, instead of using the attention maps of the last layer, we use the attention maps of the attentional pooler. We found this variant to work better.
Raw Attention: Similarly, the Raw Attention baseline was adjusted by substituting the attention maps from the last self-attention layer with those from the attentional pooler.
Other Baselines: For the remaining baseline methods, no alterations were necessary. These methods were inherently compatible with the attentional pooler, and thus could be applied directly without any further adaptation.
The adaptations described above ensure that each baseline method is appropriately tailored to the ViTs with attentional poolers, allowing for a fair comparison with our proposed LeGrad method.
9.4 Mitigation of sensitivity to irrelevant regions:
We observe that for all evaluated explainability methods, SigLIP displays high activations in image regions corresponding to the background. These activations appeared to be invariant to the input, regardless of the gradient computation’s basis, these regions were consistently highlighted. To address this issue, we computed the explainability map for a non-informative prompt, specifically "a photo of". We then leveraged this map to suppress the irrelevant activations.
Namely, for an activation $s$ under examination, we nullify any location where the activation exceeds a predefined threshold (set at 0.8) in the map generated for the dummy prompt. Formally, let $E^{s}$ denote the explainability map for activation $s$, and $E^{empty}$ represent the map for the prompt "a photo of". The correction procedure is defined as follows:
$E^{s}_{E^{empty}>th}=0,$ (12)
where $th=0.8$. This method effectively addresses the issue without resorting to external data regarding the image content.
10 Additional results
10.1 Image Classification
In this section, we extend our evaluation of the proposed LeGrad method to Vision Transformers (ViTs) that have been trained on the ImageNet dataset for the task of image classification. The results of this evaluation are presented in Table 7, also providing a comparison with other state-of-the-art interpretability methods.
It shows that LeGrad achieves superior performance on the perturbation-based benchmark, particularly in scenarios involving positive perturbations.
Another observation is that even elementary interpretability approaches, such as examining the raw attention maps from the final attention layer of the ViT, demonstrate competitive results. In fact, these basic methods surpass more complex ones like GradCAM (achieving an AUC of 53.1 versus 43.0 for negative perturbations).
Negative Positive
Method Predicted $\uparrow$ Target $\uparrow$ Predicted $\downarrow$ Target $\downarrow$
rollout [1] 53.10 53.10 20.06 20.06
Raw attention 45.55 45.55 24.01 24.01
GradCAM [37] 43.17 42.97 26.89 26.99
AttentionCAM [13] 41.53 42.03 33.54 34.05
Trans. Attrib. [12] 54.19 55.09 17.01 16.36
Partial-LRP [44] 50.28 50.29 19.82 19.80
CheferCAM [13] 54.68 55.70 17.30 16.75
LeGrad 54.72 56.43 15.20 14.13
Table 7: SOTA comparison on ViT-B/16: Comparison of interpretability methods on perturbation-based tasks on ImageNet-val for a ViT trained on ImageNet.
10.2 Pertubation example
Figure 8 illustrates the perturbation-based benchmark of Section 4.1 in the main paper. Given the explainability map generated by the interpretatability method, for the negative (respectively positive) we progressively remove the most important (respectively the least important) part of the image. We then look at the decline in the model accuracy.
Figure 8: Example of positive/negative perturbations: illustration of positive and negative perturbations used in the perturbation-based benchmark. (Top row): positive perturbation. Bottom row: negative perturbations
10.3 Qualitative examples
Finally, Figure 9 provides additional qualitative comparisons with state-of-the-art explainability methods, illustrating the efficacy of the proposed approach.
Figure 9: Qualitative Comparison to SOTA: visual comparison of different explainability methods on images from OpenImagesV7
11 Personal and human subjects data
We acknowledge the use of datasets such as ImageNet and OpenImagesV7, which contain images sourced from the internet, potentially without the consent of the individuals depicted. We also recognize that the VL models used in this study were trained on the LAION-2B dataset, which may include sensitive content. We emphasize the importance of ethical considerations in the use of such data.